# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v4

@@ I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], whereby proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
@@
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21]. And in a following work [22] the authors presented a $q$-ary $2$-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
@@
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29]. Schoeny’s result was generalized to non-binary alphabets in [30, 31]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[32]. It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 21, 35, 36, 32].
@@
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. We call such codes $2$-$b$-burst correcting codes. To the best of our knowledge, there are no explicit results about such codes. A related result can be found in a work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively. For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22]. Note that in a burst of deletions, all deletions occur consecutively. Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and $2$-$b$-burst-deletion correcting codes (where $b>1$). This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique? Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
@@
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
@@
  • A binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for all $b>1$;
@@
  • A $q$-ary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for all $q>2$ and $b>1$;
@@
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
@@
Here, it is assumed that $q$ and $b$ are constants with respect to $n$. Therefore, our results show that for $2$-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
@@
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV deals with codes for correcting two $b$-burst-deletions. In Section V, we give a new construction of non-binary two-deletion correcting code. Finally Section VI concludes this paper.
@@ II Preliminaries
In this section, we first introduce some necessary definitions, auxiliary conclusions and related results.
@@
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, $1$, 00 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $11\leq i\leq n-b+$ 1.
@@
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlap substrings of length $b$ from $\mathbf{x}$.
@@
By assumption, there is $11\leq i_{1}\leq n-b+$ 1 and $11\leq i_{2}\leq n-2b+$ 1 such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $11\leq i_{2}\leq i_{1}-$ 1. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
@@
[…]
@@
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct $t$-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
@@
Let $n^{\prime}$ and $n$ be two positive integers satisfying $n^{\prime}<n$. For each sequence $\mathbf{x}\in\Sigma_{q}^{n}$, let $\widetilde{\mathbf{x}}$ be the zero padding of $\mathbf{x}$ to the shortest length that is greater than $n$ and is divisible by $n^{\prime}$, that is, $\widetilde{\mathbf{x}}=\mathbf{x}0^{\left\lceil n/n^{\prime}\right\rceil n^{\prime}-n}$, and then $n^{\prime}\mid\left|\widetilde{\mathbf{x}}\right|$. We can represent $\mathbf{x}$ as an $n^{\prime}\times\left\lceil n/n^{\prime}\right\rceil$ array $A\left(\mathbf{x},n,n^{\prime}\right)=\left[a_{i,j}\right]$, where $a_{i,j}=\tilde{x}_{i+n^{\prime}j}$ for all $1\leq i\leq n^{\prime}$ and $10\leq j\leq\left\lceil n/n^{\prime}\right\rceil-$ 1. In other words, the $i$-th row of $A\left(\mathbf{x},n,n^{\prime}\right)$ is
$A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
@@
$A\left(\mathbf{x},7,2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}.$
@@
$A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$. Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
@@
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{2}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.

### v5

@@ I Introduction
Asubset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], whereby proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|\mathcal{C}\right|\right)$ [^1] All logarithms in this paper are to the base $2$., is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
@@
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting code with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary 22-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21]. And in a following work [22] the authors presented a $q$-ary 22-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
@@
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29]. Schoeny’s result was generalized to non-binary alphabets in [30, 31]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[32]. It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 35, 21, 36, 37, 32]. For readers’ convenience, we summarize previous results on codes correcting bursts of deletions in Table I.
@@
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. We call such codes 22-$b$-burst correcting codes. To the best of our knowledge, there are no explicit results about such codes. A related result can be found in a work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively ($(t,t_{L})$ burst deletions, for short). Let $t=2$ and $t_{L}=b$. Then their result gives a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively. For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22]. Note that in a burst of deletions, all deletions occur consecutively. Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and 22-$b$-burst-deletion correcting codes (where $b>1$). This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique? Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
@@
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of 22-$b$-burst-deletion correcting codes;
@@
  • A binary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for any $b>1$;
@@
  • A $q$-ary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for any $q>2$ and $b>1$;
@@
  • A new construction of $q$-ary two-deletion correcting codes of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for any $q>2$.
@@
Here, it is assumed that $q$ and $b$ are constants with respect to $n$. Therefore, our results show that for 22-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
TABLE I: Previous codes correcting bursts of deletions and corresponding methods
Redundancies of $q$-ary Codes References      Core Methods
single burst
of size
exactly $b$
$b\log(n/b+1)$
($q=2$)
[28]
1. representing each codeword as an array with $b$ rows
2. imposing a VT constraint on each row
$\log n+(b-1)\log\log n+O(1)$
($q=2$, $q>2$)
[29]
[30]
1. representing each codeword as an array with $b$ rows
2. encoding the first row with a VT code with run-length limited constraint
3. encoding the rest rows with shifted VT codes
$\log n+O(1)$ ($q\geq 2$) [32]
1. representing each codeword as an array with $b$ rows
2. representing this array as a $q^{b}$-ary vector
3. imposing two types of sum constraints on this vector
4. consider the signature of this vector
5. imposing a VT-type constraint together with three types of sum constraints
   on the signature
single burst
of size
at most $b$
$\log n+1$ ($b=2,q=2$)
[27]
[36]
1. imposing a sum constraint on rank sequences of codewords
       or
2. imposing a VT-type constraint on differential sequences of codewords
$(b-1)\log n+\left(\binom{b}{2}-1\right)\log\log n$
            +O(1)+O(1)
($q=2$)
[29]
   extension of the construction of $b$-burst-deletion correcting codes
   in the same work
$\left\lceil\log b\right\rceil\left(\log n+\binom{b+1}{2}\log\log n\right)$
           +O(1)+O(1)
($q=2$)
[33]
1. denoting each $t\in[1,b]$ as $t=2^{i}\cdot j$, where $i\geq 0$ and $j$ is odd
2. representing each code word as an array with $2^{i}$ rows
3. imposing a VT-type constraint alongside a “balanced” constraint on the
   first row to approximate the error positions
4. representing each codeword as an array with $t$ rows and encoding
   each row with a shifted VT code
$\log n+\binom{b+1}{2}\log\log n+O(1)$
($q=2$)
[35]
1. each codeword $\mathbf{x}$ is required to be $(\bm{p},\delta)$-dense
2. associating with $\mathbf{x}$ a vector of integers $a_{\bm{p}}(\mathbf{x})$
3. imposing a constraint on the number of $\bm{p}$ in $\mathbf{x}$ and a VT-type constraint
   on $a_{\bm{p}}(\mathbf{x})$ to approximate positions of errors
5. for each $1\leq t\leq b$, representing $\mathbf{x}$ as an array with $t$ rows and
   encoding each row with shifted VT codes
$4\log n+o\left(\log n\right)$ [34] syndrome compression technique
$\log n+8\log\log n+o\left(\log\log n\right)$
($q=2$)
[21]
1. applying the same method in [35] to approximate positions of errors
2. applying a code with $4\log n+o(\log n)$ redundant bits to correct burst
   deletions in short intervals
$\log n+(8\log q+8)\log\log n$
      $o\left(\log q\log\log n\right)$
($q>2$ is even)
1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
2. encoding the first row with the binary code in the same work
3. applying a code with $4\log n+o(\log n)$ redundant bits to correct
   errors in remaining rows
$\log n+8\log\log n+o\left(\log\log n\right)$
($q\geq 2$)
[37]    generalization of binary codes in [21]
$\log n+\log q\log\log n+O(1)$
($b=2$, $q>2$ is even)
[36]
1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
2. encoding the first row with the binary code in the same work
   with additional pattern-limited constraint
3. applying a $P$-bounded version of the binary code to remaining rows
$\log n+b\log\log n+O(1)$
($q\geq 2$)
[32]
1. associating each codeword with a binary sequence
2. applying the same method in [35] to the binary sequence to approximate
   error positions
3. for each $1\leq b^{\prime}\leq b$, applying a bounded $b^{\prime}$-burst-deletion correcting code
   to correct errors in short intervals
$(t,t_{L})$
burst
deletions
$4t\log n+o(\log\log n)$
($q=2$)
[34]
1. $t$-mixed sequences
2. syndrome compression technique
@@ 1. $t$-mixed sequences
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV deals with codes for correcting two $b$-burst-deletions. In Section V, we give a new construction of non-binary two-deletion correcting codes. Finally Section VI concludes this paper.
@@ II Preliminaries
In this section, we introduce some necessary definitions, auxiliary conclusions and related results.
@@
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, 11, 0 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $1\leq i\leq n-b+1$.
@@
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlapping substrings of length $b$ from $\mathbf{x}$.
@@
By assumption, there is $1\leq i_{1}\leq n-b+1$ and $1\leq i_{2}\leq n-2b+1$ such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $1\leq i_{2}\leq i_{1}-1$. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
@@
$\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\Sigma_{q}^{n-tb}:\begin{array}[]{c}\mathbf{y}\text{ is obtained from }\mathbf{x}\\ \text{by }t\;b\text{-burst-deletion(s)}\end{array}\right\}.$
@@
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct 22-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
@@
[…]
$A\left(\mathbf{x},n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$. If $n^{\prime}$ is clear from the context, we will denote $A\left(\mathbf{x},n^{\prime}\right)$ by $A(\mathbf{x})$.
@@
$A\left(\mathbf{x},2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}$.
@@
$A\left(\mathbf{x},3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}$.
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$. Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
@@
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.